OpenStudy (anonymous):

find an equation of the line having given the slope and containing the given point m=-4,(3,7) the equation of the line is y=?

OpenStudy (blockcolder):

Substitute them in y=m(x-x1)+y1. y=-4(x-3)+7=-4x+12+7=-4x+19. The required equation is y=-4x+19.